# Ricci-Notation Tensor Framework for Model-Based Approaches to Imaging

### 10 Jan 2024 v2

@@ Abstract
Model-based approaches to imaging, like specialized image enhancements in astronomy, facilitate explanations of relationships between observed inputs and computed outputs. These models may be expressed with extended matrix-vector (EMV) algebra, especially when they involve only scalars, vectors, and matrices, and with n-mode or index notations, when they involve multidimensional arrays, also called numeric tensors or, simply, tensors. While this paper features an example, inspired by exoplanet imaging, that employs tensors to reveal embedded 2D fast Fourier transforms in an image enhancement model, the work is actually about the tensor algebra and software, or tensor frameworks, available for model-based imaging. The paper proposes a Ricci-notation tensor (RT) framework, comprising a dual-variant index notation, with Einstein summation convention, and codesigned object-oriented software, called the RTToolbox for MATLAB. Extensions to Ricci notation offer novel representations for entrywise, pagewise, and broadcasting operations popular in EMV frameworks for imaging. Complementing the EMV algebra computable with MATLAB, the RTToolbox demonstrates programmatic and computational efficiency thanks to careful design of tensor and dual-variant index classes. Compared to its closest competitor, also a numeric tensor framework that uses index notation, the RT framework enables superior ways to model imaging problems and, thereby, to develop solutions.
@@ I Introduction
When literature on learning-based [1][2] and model-based [3][15] approaches to imaging use the word, “tensor” mainly or solely means a numeric data structure, usually \@iaciMDA MDA, together with numeric operations, like those identified with overloadable operators in object-oriented programming languages. This paper uses the word likewise, agnostic to requirements for deeper tensor character addressed, for example by Synge and Schild [16], in books on geometric tensor calculus.
@@
This paper proposes the Ricci-notation tensor (RT) framework for model-based approaches to imaging. A successor to Harrison and Joseph’s numeric tensor (NT) framework [27], itself a complement to the EMV framework, the RT framework inherits their advantages. The proposed framework comprises RT algebra and RT software, defined as the RTToolbox [33], contributed with this paper, plus MATLAB. The RT software parses the RT algebra, dispatching calculations to the MATLAB kernel, with no dependencies on other toolboxes.
Both the NT and proposed RT algebras are index notations that exploit and extend the Einstein summation convention. However, using a dual-variant index notation, the RT algebra proves equally expressive as the NT algebra, shown to be a multi-variant index notation, and aligns well with the Ricci notation of tensor calculus. Compared to NT algebra, the RT algebra supports additional outer operations. Inspired by the broadcasting of EMV algebras, they remain grounded in a dual-variant index notation with summation convention.
@@
An image enhancement example, to illustrate the RT algebra, is inspired by Sirbu et al.’s model-based research [36, 37, 38] for coronagraphy [39]. The example adopts a simplified model that involves two planes related by 2D discrete Fourier transforms and a phase aberration. Starting with \@iaciCC0 Creative Commons public domain (CC0) image [40], called Airy_disk_D65, the impact of the model on a synthetic ground-truth image having an occulted star with nearby exoplanets is computed. The exoplanets become visible after correction of the aberration, treated as unknown.
@@ II Tensor Algebra
In the proposed RT algebra, tensors are scalars, vectors, and matrices of any degree, where degree is the number of true-variant and/or false-variant indices. First, benefits of exploiting and extending Ricci notation for this purpose is introduced in the context of scalars, even though these benefits also apply to nonscalars. Second, additional extensions of the RT algebra are summarized in the context of matrices and vectors.
@@ II-A Dual-Variant Index Notation
The proposed RT algebra is introduced with reference to the NT algebra. Using an underline operator in what was claimed to be a single-variant index notation, the NT algebra exploited and extended Einstein summation to support arbitrary mixtures of inner, entrywise, and outer products of $N$-degree scalars, not called $N$-dimensional arrays with reason. Use of the word degree, or order, instead of dimensional, for the number of indices, is consistent with a geometric tensor calculus.
In the NT algebra, the underline operator prevents Einstein summation, resulting in entrywise products. Unnecessary complexity of the notation is demonstrable with simple examples, where indices are suppressed in inline math for readability. Consider the ternary inner product, $x$, and the ternary entrywise product, $y$, of three degree-one scalars, $a$, $b$, and $c$:
@@
For a variety of derivations, one needs to pair operands of $N$-ary products arbitrarily. This can be done with commutation and association identities, as follows for these examples:
@@
The underline operator stacks arbitrarily, meaning there is a countable infinite number of variants for any index. This freedom has implications for the NT software. It also means the NT algebra is actually a multi-variant index notation.
Setting aside questions of contructivism, there is a simpler formalism, namely the RT algebra, that enjoys the accepted advantages of the NT algebra for numeric tensor purposes. With the RT algebra, inspired by the Ricci notation of geometric tensor calculus, any index always has two variants, true and false, that are identified usually with typeset position, subscript and superscript, respectively. The covariant and contravariant indices of the Ricci notation inspired these names.
@@
The examples of ternary inner, entrywise, and outer products are convenient for illustrating the known Ricci operation of contraction and an extended Ricci operation called attraction, both of which are unary operators that reduce degree:
@@
With the overscript bar notation, one could instead write a sensible inequality to represent the assymetry of a tensor:
@@
Mixtures of inner, entrywise, and outer products, and ways in which to rearrange them, are significant to CP derivations like those used to illustrate the NT algebra. Consider the following key step of a degree-three CP derivation, which uses the RT overscript bar notation to clarify index position and inner parentheses to group for asymptotic efficiency:
@@
As in the EMV algebra, a left (or right) division implies the solution to a linear system. In this case, it is as follows:
@@ II-B Nonscalars and Broadcasting
Here, first and second indices of scalar operands correspond to rows and columns, respectively, of nonscalar operands. For an $N$-ary product of any-degree scalars, the operand sequence may be rearranged arbitrarily. This is not possible, in general, for an equivalent $N$-ary product of degree-zero matrices.
@@
Outer RT operations resemble broadcasting, popularized by the implied EMV algebra of Python+NumPy, but are modelled with an index notation. Variants aside, one advantage of index-based outer operations is that they express broadcasting in a way that would look familiar to high-school students. This attests further to the value of the RT algebra for model-based approaches, which favour human intelligence, as opposed to learning-based approaches, which favour machine intelligence. By design, the RT algebra includes and extends, with outer RT operations, the implied EMV algebra of MATLAB.
@@
$\alpha_{\ell}^{j}$$\equiv\begin{cases}1,&\ell=j\text{,}\\ 0,&\text{otherwise}\text{,}\end{cases}$ (48)
$\beta_{\ell}^{j}$$\equiv\begin{cases}1,&\ell=J_{1}+j\text{,}\\ 0,&\text{otherwise}\text{.}\end{cases}$ (49)
@@ III Imaging Example
After setting up a problem, inspired by an exoplanet imaging instrument called a coronagraph, this section applies the RT algebra and MATLAB to solve it. Results are presented first to summarize without equations or code a scalar optimization, called the phase aberration correction. Details, like a gradient matrix, are then formulated using the RT algebra and (inverse) DFT operators. Leveraging (inverse) fast Fourier transforms, asymptotically-efficient code is lastly presented.
@@ III-A Simulation Results
Figure 1 introduces a model-based example. It displays in false color, after each pixel value is squared, a ground-truth image synthesized from the grayscale version of \@iaciCC0 CC0-licensed source image, Airy_disk_D65. The latter simulates a focused spot of white light passing through a circular aperture and lens. Rings in the focal or image plane arise due to diffraction. The figure includes, with circles overlaid to define an annular mask, the source image in false color. Using the mask to model a coronagraph, a bright spot is blacked out, representing an occulted star. An outer symmetric region, where diffraction rings are faint, is likewise masked.
@@
Considering coronagraphy, an entrywise square root was applied in making the ground-truth from the source, after conversion from class uint8 to double and division by 255. Both images have $401\times 401$ pixels. Filled circles introduced into the annular foreground simulate twin exoplanets in the ground-truth. Using bicubic and nearest-neighbour interpolation, respectively, ground-truth and mask images are downsampled or upsampled as needed, to vary dimension sizes equally for simulation purposes. The mask is of class logical.
The ground-truth image does not exhibit a phase aberration. To synthesize an aberrated image, a random phase offset is introduced entrywise in the Fourier domain, often called the pupil plane in the context of exoplanet imaging. Because pixel values in the image plane, the inverse Fourier domain, must be real, this phase aberration is constrained to obey the required symmetry. Notwithstanding a few entries, such as the DC value, each entry of the random offset, an image-like matrix, belongs to a uniform distribution of $-\pi$ to $+\pi$ radians.
@@
In practice, a multistep computing problem is solveable only if it requires a reasonable amount of processing time, measurable in seconds ($\mathrm{s}$), and a reasonable amount of memory space, measurable in bytes ($\mathrm{B}$) of random access memory (RAM), to solve it. For desktop computing, reasonable space means the RAM of \@iaciPC personal computer (PC) suffices for the image resolution, $401\times 401$ pixels, at hand. Reasonable time means a fractional second per step suffices on the PC at hand, a Dell Latitude E7450 with $8\,\mathrm{GB}$ of RAM, for the same.
An accurate and efficient SSE software function is required for dimension sizes of interest, given a phase aberration, plus the same for first and second SSE derivatives with respect to the 2D aberration. Otherwise, the optimization may fail, due to numerical errors in approximations of the first derivative, the gradient, or may not succeed in reasonable time, due to higher-order curvature not captured without a second derivative, the Hessian, or second-derivative information, \@iaciHMF Hessian multiply function (HMF). The RT algebra enables accurate and efficient SSE, gradient, and HMF functions in MATLAB.
@@ III-B Scalar Error
The image enhancement input is an $M\times N$ aberrated image, $\mathbf{X}^{\text{a}}$. One defines a scalar SSE function in the image plane, an inverse Fourier domain, with a model. Corresponding to \@iaciSSE SSE minimum, the image enhancement output is an $M\times N$ corrected image, $\mathbf{X}^{\text{t}}$, an estimate of the ground-truth. Figure 2 shows the images, $\mathbf{X}^{\text{a}}$ and $\mathbf{X}^{\text{t}}$, after entrywise squares.
@@
By the problem definition, the aberrated image, $\mathbf{X}^{\text{a}}$, is real, and so is the corrected image, $\mathbf{X}^{\text{t}}$. Thus, the complex-valued 2D DFT, $\mathbf{Y}^{\text{a}}$, exhibits a phase symmetry. The 2D DFT of the corrected image, $\mathbf{Y}^{\text{t}}$, differs only in phase due to an offset. Therefore, for the corrected image to be real, there is an implied symmetry contraint on the phase aberration, $\mathbf{\Phi}$. Discarding the imaginary part, after an inverse 2D DFT, deals with round-off error issues that may arise in practice:
@@ III-D Hessian Function
Asymptotically, it takes equal time and space to construct the SSE and gradient as the SSE alone. Using explicit DFT operators, either takes $O(M^{2}N+MN^{2})$ time and $O(MN)$ space. Each DFT operator represents a Fourier transform, or its inverse, as matrix multiplication. One need not construct the operator, a useful formalism, as the FFT constructs \@iaciDFT DFT faster. Using 2D FFTs, the SSE and gradient or the SSE only takes $O(MN\log MN)$ time and $O(MN)$ space.
Non-asymptotically, it takes longer to compute the SSE and gradient than it does to compute the SSE alone. Yet without the gradient, on small-format test cases, the trust-region algorithm was too slow. For medium and large-format test cases, even the benefit of an explicit first-order partial derivative, the gradient, was insufficient. For a practical solution, an implicit second-order partial derivative, \@iaciHMF HMF, was needed. The RT algebra helps to formulate and simplify the desired HMF.
@@
On the other hand, the HMF, $\mathbf{F}$, is a degree-one matrix that requires at least $O(MNP)$ space and time to construct, where the dimension size, $P$, refers to a third index, $k$, of a phase step, $\Delta\phi$, a degree-three, not a degree-two, scalar here:
@@
$\Delta\mathbf{X}_{k}^{\text{t}}$ =$=\left\{\frac{\jmath}{MN}\bar{\mathbf{U}}^{\mathrm{T}}(\mathbf{Y}^{\text{t}}\circ\mathbf{e}_{i}\mathbf{e}_{j}^{\mathrm{T}})\bar{\mathbf{V}}\right\}\Delta\phi_{k}^{ij}\text{,}$ (102)
@@ III-E Implementation Details
Figure 5: Model-based SSE and gradient. Developed for the imaging application, this MATLAB function’s output arguments are the SSE, $E$, the SSE gradient, $\mathbf{\nabla}E$, and a corrected image, $\mathbf{X}^{\text{t}}$. Its input arguments are a phase aberration, $\mathbf{\Phi}$, the aberrated image, $\mathbf{X}^{\text{a}}$, and the occultation (background) mask, $\mathbf{W}^{\text{b}}$.
@@
The MATLAB functions, ssefun and hessmfun, do not require the RTToolbox. However, they exploit efficiencies derived using the RT algebra. Figures 5 and 6 clarify the 2D (inverse) DFTs revealed after algebraic manipulations that involved purely numeric aspects of tensor operands, some nonzero degree. The dual-variant index notation excelled on a variety of inner, entrywise, outer, and mixed products.

### 08 Apr 2024 v3

@@ Abstract
Model-based approaches to imaging, like specialized image enhancements in astronomy, facilitate explanations of relationships between observed inputs and computed outputs. These models may be expressed with extended matrix-vector (EMV) algebra, especially when they involve only scalars, vectors, and matrices, and with n-mode or index notations, when they involve multidimensional arrays, also called numeric tensors or, simply, tensors. While this paper features an example, inspired by exoplanet imaging, that employs tensors to reveal (inverse) 2D fast Fourier transforms in an image enhancement model, the work is actually about the tensor algebra and software, or tensor frameworks, available for model-based imaging. The paper proposes a Ricci-notation tensor (RT) framework, comprising a dual-variant index notation, with Einstein summation convention, and codesigned object-oriented software, called the RTToolbox for MATLAB. Extensions to Ricci notation offer novel representations for entrywise, pagewise, and broadcasting operations popular in EMV frameworks for imaging. Complementing the EMV algebra computable with MATLAB, the RTToolbox demonstrates programmatic and computational efficiency via careful design of numeric tensor and dual-variant index classes. Compared to its closest competitor, also a numeric tensor framework that uses index notation, the RT framework enables superior ways to model imaging problems and, thereby, to develop solutions.
@@ I Introduction
When literature on learning-based [1][2] and model-based [3][15] approaches to imaging use the word, “tensor” mainly or solely means a numeric data structure, usually an MDA, together with numeric operations, like those identified with overloadable operators in object-oriented programming languages. This paper uses the word likewise, avoiding requirements for deeper tensor character addressed, for example by Synge and Schild [16], in books on geometric tensor calculus.
@@
This paper proposes the Ricci-notation tensor (RT) framework for model-based approaches to imaging. A successor to Harrison and Joseph’s numeric tensor (NT) framework [27], itself a complement to an EMV framework, the RT framework inherits their advantages. The proposed framework comprises RT algebra and RT software, defined as the RTToolbox [33], developed with this paper, plus MATLAB. The RT software parses the RT algebra, dispatching calculations to the MATLAB kernel, with no dependencies on other toolboxes.
Both the NT and proposed RT algebras are index notations that exploit and extend the Einstein summation convention. However, using a simpler dual-variant index notation, the RT algebra proves equally expressive as the NT algebra, shown here to be a multi-variant index notation, and aligns well with the Ricci notation of tensor calculus. Moreover, the RT algebra supports additional outer operations. Inspired by the broadcasting of EMV algebras, they remain grounded in a dual-variant index notation with summation convention.
@@
An image enhancement example, developed here, illustrates the RT algebra. Inspired by Sirbu et al.’s model-based research [36, 37, 38] for coronagraphy [39], the example adopts a simple coronagraph model with two planes related by 2D discrete Fourier transforms and a phase aberration. Starting with a Creative Commons public domain (CC0) image [40], called Airy_disk_D65, the impact of the model on a synthetic ground-truth image having an occulted star with nearby exoplanets is computed. The exoplanets become visible after correction of the aberration, treated as unknown.
@@ II Tensor Algebra
In the proposed RT algebra, tensors are scalars, vectors, and matrices of any degree, where degree is the number of true-variant and/or false-variant indices. First, benefits of exploiting and extending Ricci notation for this purpose are introduced in the context of scalars, even though these benefits also apply to nonscalars. Second, additional extensions of the RT algebra are summarized in the context of vectors and matrices.
@@ II-A Dual-Variant Index Notation
The proposed RT algebra is introduced with reference to the NT algebra. Using an underline operator in what its authors called a single-type index notation, the NT algebra exploited and extended Einstein summation to support arbitrary mixtures of inner, entrywise, and outer products of $N$-degree scalars, not called $N$-dimensional arrays with reason. Use of the word degree (or order), instead of dimensional, for the number of indices is consistent with a geometric tensor calculus.
In the NT algebra, the underline operator prevents Einstein summation, yielding entrywise products. One can demonstrate unnecessary complexity of the notation with simple examples. Consider the ternary inner product, $x$, and the ternary entrywise product, $y$, of three degree-one scalars, $a$, $b$, and $c$, with indices suppressed in inline math for readability:
@@
For a variety of derivations, one needs to pair operands of $N$-ary products arbitrarily. This is done with commutation and association identities, as follows for these examples:
@@
The underline operator stacks arbitrarily, meaning any index has a countable infinite number of variants. This freedom has implications for the NT software. It also means the NT algebra is actually a multi-type or multi-variant index notation.
Setting aside questions of contructivism, consider a simpler formalism, namely the RT algebra, that enjoys the accepted advantages of the NT algebra for numeric tensor purposes. With the RT algebra, inspired by the Ricci notation of geometric tensor calculus, any index always has two variants, true and false, that are identified usually with typeset position, subscript and superscript, respectively. The covariant and contravariant indices of the Ricci notation motivated these names.
@@
The examples of ternary inner, entrywise, and outer products are convenient for illustrating the known Ricci operation of contraction and an extended Ricci operation called attraction, both of which are unary operations that reduce degree:
@@
With the overscript bar notation, one could instead write a sensible inequality to represent the asymmetry of a tensor:
@@
Mixtures of inner, entrywise, and outer products, and ways in which to rearrange them, are significant to CP derivations like those used to illustrate the NT algebra. Consider the following key step of a degree-three CP derivation, which uses the RT overscript bar notation to clarify index positions and inner parentheses to group for asymptotic efficiency:
@@
As with EMV algebras, a left (or right) division implies the solution to a linear system. In this case, it is as follows:
@@ II-B Nonscalars and Broadcasting
Here, first and second indices of scalar operands correspond to rows and columns, respectively, of nonscalar operands. For an $N$-ary product of degree-two scalars, the operand sequence may be rearranged arbitrarily. This is not possible, in general, for an equivalent $N$-ary product of degree-zero matrices.
@@
Outer RT operations resemble broadcasting, popularized by the implied EMV algebra of Python+NumPy, but are modelled with an index notation. Variants aside, one advantage of index-based outer operations is that they express broadcasting in a way that would look familiar to high-school students. This attests to the special value of the RT algebra for model-based approaches, which favour human intelligence, in relation to learning-based approaches, which favour machine intelligence. By design, the RT algebra includes and extends, with outer RT operations, the implied EMV algebra of MATLAB.
@@
$\alpha_{\ell}^{j}$ =$=\begin{cases}1,&\ell=j\text{,}\\ 0,&\text{otherwise}\text{,}\end{cases}$ (48)
$\beta_{\ell}^{j}$ =$=\begin{cases}1,&\ell=J_{1}+j\text{,}\\ 0,&\text{otherwise}\text{.}\end{cases}$ (49)
@@ III Imaging Example
After proposing a problem, inspired by an exoplanet imaging instrument called a coronagraph, this section applies the RT algebra and MATLAB to solve it. Results, presented first, summarize without equations or code a scalar optimization, called the phase aberration correction. Details, like a gradient matrix, are then modeled and simplified using (inverse) DFT operators and the RT algebra. Code, presented last, leverages (inverse) fast Fourier transforms for efficiency.
@@ III-A Simulation Results
Figure 1 introduces a model-based example. It displays in false color, after each pixel value is squared, a ground-truth image synthesized from the grayscale version of a CC0-licensed source image, Airy_disk_D65. The latter simulates a focused spot of white light passing through a circular aperture and lens. Rings in the focal or image plane arise due to diffraction. The figure includes, with circles overlaid to define an annular mask, the source image in false color. Using the mask to model a coronagraph, a bright spot is blacked out, representing an occulted star. An outer symmetric region, where diffraction rings are faint, is likewise masked.
@@
Given coronagraphy, the example uses an entrywise square root to make the ground-truth from the source, after conversion from class uint8 to double and division by 255. Both images have $401\times 401$ pixels. Filled circles introduced into the annular foreground simulate twin exoplanets in the ground-truth. Using bicubic and nearest-neighbour interpolation, respectively, ground-truth and mask images are downsampled or upsampled as needed, to vary dimension sizes equally for simulation purposes. The mask is of class logical.
The ground-truth image does not exhibit a phase aberration. To synthesize an aberrated image, a random phase offset is introduced entrywise in the Fourier domain, often called the pupil plane in the context of coronagraphy. Because pixel values in the image plane, the inverse Fourier domain, must be real, this phase aberration is constrained to obey the required symmetry. Notwithstanding a few entries, such as the DC value, each entry of the random offset, an image-like matrix, belongs to a uniform distribution of $-\pi$ to $+\pi$ radians.
@@
In practice, a multistep computing problem is solveable only if it requires a reasonable amount of processing time, measurable in seconds ($\mathrm{s}$), and a reasonable amount of memory space, measurable in bytes ($\mathrm{B}$) of random access memory (RAM), for all steps. With desktop computing, reasonable space means the RAM of a personal computer (PC) suffices for the image resolution, $401\times 401$ pixels, at hand. Reasonable time means a fractional second per step suffices on the PC at hand, a Dell Latitude E7450 with $8\,\mathrm{GB}$ of RAM, for the same.
An accurate and efficient SSE software function is required for dimension sizes of interest, given a phase aberration, plus the same for first and second SSE derivatives with respect to the 2D aberration. Otherwise, the optimization may fail, due to numerical errors in approximations of the first derivative, the gradient, or may not succeed in reasonable time, due to higher-order curvature not captured without a second derivative, the Hessian, or second-derivative information, a Hessian multiply function (HMF). The RT algebra enables accurate and efficient SSE, gradient, and HMF functions in MATLAB.
@@ III-B Scalar Error
The image enhancement input is an $M\times N$ aberrated image, $\mathbf{X}^{\text{a}}$. One defines a scalar SSE function in the image plane, an inverse Fourier domain, with a model. Corresponding to an SSE minimum, the image enhancement output is an $M\times N$ corrected image, $\mathbf{X}^{\text{t}}$, an estimate of the ground-truth. Figure 2 shows the images, $\mathbf{X}^{\text{a}}$ and $\mathbf{X}^{\text{t}}$, after entrywise squares.
@@
By the problem definition, the aberrated image, $\mathbf{X}^{\text{a}}$, is real, and so is the corrected image, $\mathbf{X}^{\text{t}}$. Thus, the complex-valued 2D DFT, $\mathbf{Y}^{\text{a}}$, exhibits a phase symmetry. The 2D DFT of the corrected image, $\mathbf{Y}^{\text{t}}$, differs only in phase due to an offset. Therefore, for the corrected image to be real, there is an implied symmetry contraint on the phase aberration, $\mathbf{\Phi}$. Discarding the imaginary part, after an inverse 2D DFT, deals also with round-off error issues that arise in practice:
@@ III-D Hessian Function
Asymptotically, it takes equal time and space to construct the SSE and gradient as the SSE alone. Using explicit DFT operators, either takes $O(M^{2}N+MN^{2})$ time and $O(MN)$ space. Each DFT operator represents a Fourier transform, or its inverse, as matrix multiplication. One need not construct the operator, a useful formalism, as the FFT constructs a DFT faster. Using 2D FFTs, the SSE and gradient or the SSE only takes $O(MN\log MN)$ time and $O(MN)$ space.
Non-asymptotically, it takes longer to compute the SSE and gradient than it does to compute the SSE alone. Yet without the gradient, on small-format test cases, the trust-region algorithm was too slow. For medium and large-format test cases, even the benefit of an explicit first-order partial derivative, the gradient, was insufficient. For a practical solution, an implicit second-order partial derivative, an HMF, was needed. The RT algebra helps to formulate and simplify the desired HMF.
@@
On the other hand, the HMF, $\mathbf{F}$, is a degree-one matrix that requires at least $O(MNP)$ space and time to construct, where one dimension size, $P$, refers to a third index, $k$, of a phase step, $\Delta\phi$, a degree-three, not a degree-two, scalar here:
@@
$\Delta\mathbf{X}_{k}^{\text{t}}$ $=\frac{\jmath}{MN}\bar{\mathbf{U}}^{\mathrm{T}}(\mathbf{Y}^{\text{t}}\circ\mathbf{e}_{i}\mathbf{e}_{j}^{\mathrm{T}})\bar{\mathbf{V}}\,\Delta\phi_{k}^{ij}\text{,}$ (102)
@@ III-E Implementation
Figure 5: Model-based SSE and gradient. Developed for the imaging example, this MATLAB function’s output arguments are the SSE, $E$, the SSE gradient, $\mathbf{\nabla}E$, and a corrected image, $\mathbf{X}^{\text{t}}$. Its input arguments are a phase aberration, $\mathbf{\Phi}$, the aberrated image, $\mathbf{X}^{\text{a}}$, and the occultation (background) mask, $\mathbf{W}^{\text{b}}$.
@@
The MATLAB functions, ssefun and hessmfun, do not require the RTToolbox. However, they exploit efficiencies derived using the RT algebra. Figures 5 and 6 clarify the (inverse) 2D DFTs revealed after algebraic manipulations that involved purely numeric aspects of tensor operands, some nonzero degree. The dual-variant index notation excelled on a variety of inner, entrywise, outer, and mixed products.